Marrie is half age of Ann when Ann was 6 years old and and now Ann is 34 years old how old is marrie?

Answer by stanbon(75887)

Marrie is half age of Ann when Ann was 6 years old and and now Ann is 34 years old how old is marrie?
-------
Equations:
M = (1/2)A when A = 6
Then M was 3.
So, Ann is 3 years older than Marrie.
----
Ans: If Ann is 34, Marie is 34-3 = 31
-------------
Cheers,
Stan H.
--------------